OpenStudy (anonymous):

Ok, how about this one:) 4(x-1) = 6(x-15). Could someone help me with this one? Denise

OpenStudy (anonymous):

the answer should be 43, I just need to know how to get that;)

OpenStudy (anonymous):

4x-4=6x-90 2x=86 x=43

OpenStudy (anonymous):

Ok, distributive property right, I got it, I think:) Thank you much smartwish:) for taking the time, denise

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

wait, how did you get the 2x?

OpenStudy (anonymous):

Subtracting 4x from both sides 4x-4x=0 6x-4x=2x

OpenStudy (anonymous):

Ok, thank you again!!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

shoot, I get a -43:(

OpenStudy (anonymous):

4x-4=6x-90 add 90 to both sides

OpenStudy (anonymous):

Ok, I'll work with it til I understand, you have given me all the details I should need:) thank you again:)